X Company makes two products, A and B, and uses an activity-based costing overhead allocation system, with three cost pools and three cost drivers. Budgeted costs and driver information for 2020 were as follows:

Cost Pool 1 Cost Pool 2 Cost Pool 3
Budgeted costs $104,000     $57,000     $24,000    
Cost Drivers
  Product A 80,000     55,100     75,000    
  Product B 76,100     48,800     48,100    


What will the allocation to Product B be? [round overhead rate(s) to two decimal places].
Answer #1

d

Cost Pool Budgeted Cost Cost Driver Cost Drivers B Total Cost Drivers Activity Rate ( Budgeted cost / Total cost drivers Cost allocated to Product B (Cost Drivers B x Activity Rate)
1                 104,000                    80,000       76,100       156,100 104,000/156,100 = 0.67 76,100 x 0.67= 50,700.83
2                   57,000                    55,100       48,800       103,900 57,000/103,900 = 0.55 48,800 x 0.55 = 26,771.90
3                   24,000                    75,000       48,100       123,100 24,000/123,100 = 0.19 48,100 x 0.19 = 9,377.74
$86,850

Cost allocated to product B is $86,850.

Kindly comment if you need further assistance. Thanks‼!

Note: Exact answer may slightly differ due to rounding off